A person is biking and rode 30 miles in 2 hours with the wind and rode 12 miles in 3 hours against the wind. What is the rate of the wind speed? How would I set up the problem with 2 variables?
Answer by scott8148(5891)

let b="bike speed(no wind)", and w="wind speed"

b+w=30/2 and b-w=12/3

subtracting one equation from the other will give an equation for w